John does 10 pushups on the first day of a 30-day month, and then increases the number of pushups by 3 pushups a day. How many pushups has he done in total by the end of the month?
Answer by tiffany222(56)

To find sum of an arithmetic sequence use this formula:
Sn = n/2 (a1 + an)
Known information:
n = 30, a1 = 10, d = 3
To find an (last term of sequence), use this formula:
an = a1 + (n-1)d
a30 = 10 + (30-1)(3)
a30 = 10 +87
a30 = 97
Substitute into sum formula:
Sn = n/2 (a1 + an)
S30 = 30/2 (10 + 97)
S30 = 15(107)
S30 = 1605
Final Answer: 1605 pushups